Wright Lighting Fixtures forecasts its sales in units for the next four months as follows:

March 26,000
April 28,000
May 25,500
June 24,000

Wright maintains an ending inventory for each month in the amount of three times the expected sales in the following month. The ending inventory for February (March’s beginning inventory) reflects this policy. Materials cost $8 per unit and are paid for in the month after production. Labor cost is $12 per unit and is paid for in the month incurred. Fixed overhead is $22,000 per month. Dividends of $22,000 are to be paid in May. The firm produced 25,000 units in February.

Complete a production schedule and a summary of cash payments for March, April, and May. Remember that production in any one month is equal to sales plus desired ending inventory minus beginning inventory.

Wright Lighting Fixtures
Production Schedule
March April May June
Projected unit sales
Desired ending inventory
Total units required 0 0 0
Beginning inventory
Units to be produced 0 0 0
Cash Payments
February March April May
Units produced
Material cost
Labor cost
Fixed overhead
Dividends
Total cash payments $0 $0 $0
Answer #1

Wright Lighting fixtures

Production Schedule:

Particulars March April May June
Projected Unit sales (given) (Refer Note 1) 26,000 28,000 25,500 24,000
Desired Ending inventory (Refer Note 2) 84,000 76500 72000
Total units required (refer Note 3) 110,000 104,500 97,500 24,000
Beginning inventory (refer note 4) 78000 84000 76500
Units to be produced (refer note 5) 32,000 20,500 21,000

Cash Payments:

Particulars Feb March April May
Units produced 25,000 32,000 20,500 21,000
Material cost (refer note 6) 200,000 256,000 164,000
Labor cost (refer note 7) 384,000 246,000 252,000
Fixed Overhead (refer note 8) 22,000 22,000 22,000
Dividends (refer note 9) 22,000
Total cash Payments (refer note 10) 606,000 524,000 460,000

Working Notes:

1. Sales unit forecast for the months are given in the question.

2. Ending inventory is stated as 3 times the sales forecast of the next month.

So ending inventory :

March: Sales of April * 3 = 28000 *3 = 84000

April: Sales of May *3 = 25500 *3 = 76500

May: sales of June * 3 = 24000 *3 = 720003.

3. Total units required = Sales + Ending inventory

March = 26000 + 84000 = 110000 units

April = 28000 + 76500 = 104500 units

May = 25500 + 72000 = 97500 units

4. Opening inventory :
March: Ending inventory of Feb becomes opening inventory of March. Ending inventory of Feb = Sales of March * 3

= 26000 *3 = 78000.

April: Ending inventory of March as already calculated in WN 2 above becomes opening inventory of April = 84000

May: Ending inventory of April as already calculated in WN 2 above becomes opening inventory of May = 76500.

5. Units to be produced = Total units required - beginning inventory

March: 110000 -78000= 32000

April : 104500 - 84000 = 20500

May : 97500 - 76500 = 21000

6. Material cost is paid in the next month of production. It is obtained by the formula: Material cost per unit * units produced in previous month

March : Feb production * $ 8 :  25000 units * 8 = $ 200,000

April: March production units * $8 =: 32000 * 8 = $ 256,000

May : April production units * $ 8 = 20500 * 8 = $164,000

7. Labor cost is paid in the same month when incurred. Labor cost = Units produced * $12

March: 32000 * 12 = $384,000

April: 20500 * $12 = $ 246,000

May : 21000 * $12 = $252,000

8. Fixed overhead is constant at $22,000 in all months.

9. Dividends of $22,000 is paid in May.

10. Total cash Payments: Material cost + Labor cost + Fixed overhead + Dividends

March = 200,000 + 384,000 + 22,000 = $606,000

April = 256,000 + 246,000 + 22,000 = $ 524,000

May = 164,000 + 252,000 + 22,000 + 22,000 = $460,000